8x^2-2x-3=0 solve for x. write both solutions, separated by a comma.
Answer by sofiyac(983)

(4x-3)(2x+1)=0
4x-3=0
4x=3
x=3/4
2x+1=0
2x=-1
x=-1/2
3/4,-1/2